### v1

@@ Abstract
Large language models (LLMs) have revolutionized natural language processing. However, effectively incorporating complex and potentially noisy user interaction data remains a challenge. To address this, we propose User-LLM, a novel framework that leverages user embeddings to contextualize LLMs. These embeddings, distilled from diverse user interactions using self-supervised pretraining, capture latent user preferences and their evolution over time. We integrate these user embeddings with LLMs through cross-attention and soft-prompting, enabling LLMs to dynamically adapt to user context. Our comprehensive experiments on MovieLens, Amazon Review, and Google Local Review datasets demonstrate significant performance gains across various tasks. Notably, our approach outperforms text-prompt-based contextualization on long sequence tasks and tasks that require deep user understanding while being computationally efficient. We further incorporate Perceiver layers to streamline the integration between user encoders and LLMs, reducing computational demands.
@@ 1 Introduction
Large language models (LLMs) have revolutionized the field of natural language processing (NLP) Brown et al. (2020); Chowdhery et al. (2023); OpenAI (2023); Touvron et al. (2023); Anil et al. (2023); Google (2023). With their ability to learn and adapt from massive amounts of textual data, LLMs offer significant opportunities for user modeling and personalization. By analyzing user interactions and understanding user preferences, LLMs can be leveraged to power recommendations Liu et al. (2023b); Lyu et al. (2023); Ji et al. (2023), language generation, summarization Liu et al. (2023d); Basyal & Sanghvi (2023), and question answering Kojima et al. (2022); Wei et al. (2023) in ways that are highly relevant and engaging to users. This paves the way for creating more personalized and context-aware language-based applications and services.
@@
Figure 1: Illustration of two different mechanisms to incorporate user interaction history data to LLMs. (1) User history data are expressed in natural language and fed to LLMs via text prompt; (2) User embeddings, distilled from user history data, are used to contextualize LLMs.
User interactions represent a rich source of behavioral data generated from a user’s engagement with digital systems. Spanning a wide range from textual input, search queries, media consumption (e.g., videos watched or rated), to social media activities, navigation patterns, location visits, and more, these interactions hold valuable insights for user modeling. One straightforward approach to leveraging this data with LLMs is to directly finetune LLMs on textual components, using a user’s interaction history as the text prompt. However, user interaction data is often complex, spanning multiple journeys with sparse data points, various interaction types (multimodal), and potential noise or inconsistencies. This complexity can hinder an LLM’s ability to identify and focus on the most relevant patterns. Moreover, effective personalization often requires a deep understanding of the context and latent intent behind user actions, which can pose difficulties for LLMs trained predominantly on vast, surface-level language corpora. Additionally, user interaction data, such as extended histories, can be very lengthy. Processing and modeling such long sequences (e.g., a year’s worth of history) with LLMs can strain computational resources, making it practically infeasible. Addressing these challenges is key to unlocking the full potential of LLMs in user modeling and personalization.
To address the inherent complexities and limitations of leveraging raw user interaction data with LLMs, we propose a novel approach centered around user embeddings, as illustrated in Fig. 1. These compressed representations, distilled from diverse and noisy user interactions, effectively capture the essence of a user’s behavioral patterns and preferences across various interaction modalities. By contextualizing the LLM with user embeddings during finetuning or inference, we aim to
  • Enhance its ability to identify relevant patterns navigating complexity and noise,
  • Facilitate understanding and adaptation to the latent intent, dynamic context, and temporal evolution behind user actions,
  • Mitigate the computational demands of processing extensive interaction histories by working with condensed representations.
This approach empowers LLMs with a deeper understanding of users’ historical patterns and latent intent, enabling LLMs to tailor responses and generate personalized outcomes.
Our approach consists of two key phases: generating high-quality user embeddings and contextualizing LLMs with these user embeddings. In phase one, we pretrain a Transformer-based encoder on user interaction data, utilizing self-supervised learning to capture behavioral patterns across multiple interaction modalities. We use a multi-feature autoregressive Transformer to generate embeddings that capture long-range dependencies and contextual relationships within sequential data while handling multimodal user data effectively. In phase two, we integrate user embeddings with an LLM during finetuning using cross-attention, where the LLM’s intermediate text representations attend to the output embeddings from the pretrained user encoder, enabling dynamic context injection (similar to Flamingo Alayrac et al. (2022)).
We present a comprehensive study across three public datasets and various tasks. Our findings conclusively demonstrate that the embedding-based approach successfully mitigates the limitations associated with directly finetuning LLMs on raw textual user interaction data. The approach shows great potential to enhance performance in several application domains, including user understanding, personalized recommendations, and text generation.
@@
  • We introduce User-LLM, a versatile framework that leverages user embeddings distilled from diverse modalities (e.g., video watch history, ratings, location visits) to contextualize LLMs. It dynamically incorporates user preferences and behaviors from various interaction modalities, enhancing LLM understanding and personalization capabilities while supporting various encoder architectures and multimodal fusion mechanisms.
  • We rigorously assess User-LLM against state-of-the-art baselines, demonstrating significant performance improvements, particularly in tasks requiring deep user understanding and long context inputs. Our approach excels in personalization while maintaining computational efficiency, outperforming traditional text-prompt-based contextualization.
  • User-LLM offers flexible training strategies. Notably, finetuning only the user encoder while keeping the LLM frozen effectively contextualizes the LLM and outperforms LoRA-based text-prompt LLM tuning (which also preserves original LLM weights). This highlights User-LLM’s adaptability, effectively contextualizing the LLM without altering its core, maintaining language modeling and generalization abilities.
  • We provide in-depth analyses of cross-attention and soft-prompt approaches for integrating user embeddings, exploring the potential of Perceiver layers for further efficiency gain. Extensive ablation studies offer practical guidance on the impact of different User-LLM components and design choices.
@@ 2.1 User Modeling
Dual encoders and self-supervised learning are widely employed for user models and recommendation systems, providing transferable representations as user understanding and personalization signals Covington et al. (2016b; a); Sountsov & Sarawagi (2016); Volkovs et al. (2017); Chidambaram et al. (2018); Gillick et al. (2018); Yanga et al. (2018); Ma et al. (2018); Yi et al. (2019); Gillick et al. (2019); Yang et al. (2020); Jiang et al. (2020); Ning et al. (2021). Other self-supervised learning approaches such as contrastive learning Yao et al. (2021); Xie et al. (2022); Chen et al. (2022b); Xia et al. (2023) and graph representation learning Cai et al. (2023); Yang et al. (2023) have proven to learn high-quality representations from massive but noisy user interaction data. Recently, pre-trained BERT model Devlin et al. (2018) has been applied to sequential recommendation systems. For instance, Bert4Rec Sun et al. (2019) and U-Bert Qiu et al. (2021) leverage pre-training and fine-tuning approaches to model users’ behavior in a self-supervised way. Our work explores the potential of integrating user representations with the emerging capabilities of LLMs to further improve user modeling and personalization.
2.2 Language Model based Personalization
The emergence of LLMs has propelled generative models to demonstrate superior natural language generation capabilities compared to discriminative models. There has been extensive work on different ways to format user historical interactions as prompts and to leverage LLMs to generate recommendations across various tasks Petrov & Macdonald (2023); Kang et al. (2023); Xu et al. (2023b); Liu et al. (2023b); Lyu et al. (2023); Ji et al. (2023); Li et al. (2023); Wu et al. (2023a). However, directly incorporating a user’s entire history into an LLM’s context window is often impractical or overly expensive. In contrast to the aforementioned approaches, we propose to leverage user embeddings to contextualize LLMs.
A recent study Doddapaneni et al. (2024) explored LLM personalization using user embeddings derived from user activity text. In contrast, we utilize user activity tokens for a more efficient approach. In addition, we employ sophisticated fusion techniques and conduct a comprehensive empirical evaluation across multiple datasets and tasks to provide a thorough assessment of our method’s capabilities.
2.3 Long Context in LLMs
Long context is crucial for models to incorporate long-term user data and behavior patterns. There has been extensive work to enable long context for LLMs. One line of work is to extend context window by remapping the positional encodings of the long context to the original window size Chen et al. (2023a); Jin et al. (2024), or by exposing relevant additional contexts available to a subset of attention layers Tworkowski et al. (2023). Another direction is to distill the information in the long context to fit in the existing window by using a separate retrieval model Xu et al. (2023a) or a memory bank Wang et al. (2023) etc., or to compress the information by training special tokens Ge et al. (2023); Mu et al. (2023); Chevalier et al. (2023). Other approaches include modified attention computation Ding et al. (2023); Liu et al. (2023a); Chen et al. (2023b) and linear attention frameworks such as structured state space models Gu et al. (2023); Gu & Dao (2023); Tay et al. (2022). Our method tackles long context by using a single token to represent each user event. The extracted representations align with the LLM text embedding space. Therefore, our model is compatible with existing LLM-based techniques and applications.
2.4 Multimodal LLM
Early multimodal LLMs such as CLIP Radford et al. (2021) and ImageBind Girdhar et al. (2023) relied on aligning image and text representations. Subsequently, fusion techniques leveraging cross-attention (e.g., Flamingo Alayrac et al. (2022), Coca Yu et al. (2022)) and soft prompts (e.g., PaLI Chen et al. (2022a), Palm-E Driess et al. (2023)) emerged. Recent works, such as NextGPT Wu et al. (2023b), OneLLM Han et al. (2023), and Anymal Moon et al. (2023), explored unified frameworks for diverse input modalities. Additionally, end-to-end training of multimodal models has gained attention, as seen in Gato Reed et al. (2022) and Gemini Gemini Team Google (2023). Inspired by these advances, User-LLM focuses on efficient LLM contextualization with user understanding capabilities.
3 User-LLM
Figure 2: Overview of User-LLM. Left: Multimodal Autoregressive Transformer encoder pretraining. Right: LLM contextualization with user embeddings. Features from a user’s interaction history are encoded by the Autoregressive User Encoder and then integrated into the language model via cross-attention.
@@ 3 User-LLM
This section introduces the User-LLM framework, a two-stage approach for contextualizing LLM with user embeddings, as illustrated in Fig.2. First, a pretrained user encoder generates user embeddings. Second, these embeddings are integrated into the LLM through cross-attention or soft-prompt techniques, providing the LLM additional context and guidance to generate personalized responses.
3.1 Embedding Generation
User-LLM leverages a Transformer-based encoder to generate user embeddings from ID-based feature sequences. Consider a scenario where user data contains two modalities, $a$ and $b$. Each modality has its own vocabulary. Items from each modality are mapped to integer IDs and have distinct embedding representations. To create a combined representation, aligned items $a_{i}$ and $b_{i}$ are fused into a single embedding $f_{i}$. The sequence of fused embeddings serves as input to the user encoder for generating user embeddings.
Autoregressive Encoder In this work, we employ an Autoregressive Transformer as the user encoder. As illustrated in Fig. 2 (left), this model takes a sequence of user activities, where each activity is described by multiple modalities (e.g., item name, rating, category). Individual features are embedded separately, then concatenated and processed by a Transformer decoder. The output is projected back to the original feature spaces, where modality-specific softmax layers calculate logits. Note that input embeddings and output softmax layers share weights.
The autoregressive design enables the model to predict upcoming tokens based on previous ones, using cross-entropy loss for optimization. Given a sequence of $N$ input tokens, the Autoregressive Transformer generates $N$ corresponding embeddings. These embeddings serve as user context for LLMs, enabling personalized response generation.
@@ 3.1 Embedding Generation
3.2 LLM Contextualization with User Embedding
3.2.1 Model Architecture
@@
User-LLM integrates user embeddings with LLMs using cross-attention (see Fig. 2). The output embeddings from the pretrained user encoder are cross-attended with intermediate text representations within the LLM, similar to how Flamingo Alayrac et al. (2022) works. Note that our framework remains adaptable to different integration mechanisms including prepending embeddings as soft-prompt Lester et al. (2021).
3.2.2 Efficiency
Compared to text-prompt-based methods, User-LLM offers substantial efficiency gains. First, it leverages pretrained weights of both the user encoder and LLM, improving model convergence speed with less trainable parameters. Additionally, User-LLM condenses user activities into dense representations, requiring only a single token per event. This frees up the LLM’s context window, leading to improved inference efficiency. Furthermore, User-LLM leverages cross-attention, maintaining user embeddings at a smaller dimension than the LLM’s model dimension and eliminating the need for projection operations. This design choice significantly reduces computational overhead during both training and inference.
To further optimize inference efficiency, User-LLM integrates Perceiver Jaegle et al. (2021) units into its projection layers. Perceiver is a Transformer-based architecture that utilizes a trainable latent query to extract relevant information from input representations through cross-attention. In User-LLM, Perceiver effectively compresses the user embeddings into a compact format using a latent query with shorter sequence length. This compression reduces the number of tokens needed to represent the user history, further freeing up the LLM’s context window. This, along with Perceiver’s ability to distill insights from noisy contexts, makes User-LLM ideal for practical applications with extensive user histories.
3.2.3 Training Strategies

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable success across various domains, but effectively incorporating complex and potentially noisy user timeline data into LLMs remains a challenge. Current approaches often involve translating user timelines into text descriptions before feeding them to LLMs, which can be inefficient and may not fully capture the nuances of user behavior. Inspired by how LLMs are effectively integrated with images through direct embeddings, we propose User-LLM, a novel framework that leverages user embeddings to directly contextualize LLMs with user history interactions. These embeddings, generated by a user encoder pretrained using self-supervised learning on diverse user interactions, capture latent user behaviors and interests as well as their evolution over time. We integrate these user embeddings with LLMs through cross-attention, enabling LLMs to dynamically adapt their responses based on the context of a user’s past actions and preferences.
Our approach achieves significant efficiency gains by representing user timelines directly as embeddings, leading to substantial inference speedups of up to 78.1X. Comprehensive experiments on MovieLens, Amazon Review, and Google Local Review datasets demonstrate that User-LLM outperforms text-prompt-based contextualization on tasks requiring deep user understanding, with improvements of up to 16.33%, particularly excelling on long sequences that capture subtle shifts in user behavior. Furthermore, the incorporation of Perceiver layers streamlines the integration between user encoders and LLMs, yielding additional computational savings.
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP) (Brown et al. 2020; Chowdhery et al. 2023; OpenAI 2023; Touvron et al. 2023; Anil et al. 2023; Google 2023) and demonstrated remarkable success across various domains, including language generation, summarization (Liu et al. 2023b; Basyal and Sanghvi 2023), and question answering (Kojima et al. 2022; Wei et al. 2023). This success has sparked significant interest in their potential to power the next generation of personalized AI agent - systems that understand and respond to individual needs and preferences. While LLMs have shown remarkable capabilities, their effectiveness often hinges on the quality and relevance of the data they are trained on. A key challenge lies in how to effectively incorporate the wealth of information contained within complex and potentially noisy user timeline data into these models. These timelines, spanning a wide range of interactions from textual input and search queries to media consumption, social media activity, location visits, etc., represent a rich source of user behavioral data. These data hold valuable insights crucial for deep user understanding and tailored user experience.
Existing approaches often translate user timelines into lengthy textual descriptions before feeding them into LLMs. This approach, while conceptually straightforward, leads to significantly computational overhead during inference due to the long context windows required to process these textual representation. Moreover, it may fail to capture the subtle nuances of user behavior and preferences essential for achieving truly personalized AI experiences.
@@
Figure 1: Illustration of two mechanisms to incorporate user interaction history data to LLMs. (1) Text Prompt: User history data are expressed in natural language and fed to LLMs via text prompt; (2) User-LLM: User embeddings, distilled from user history data, are used to contextualize LLMs.
Inspired by the effective integration of LLMs with other modalities, such as images (Alayrac et al. 2022; Chen et al. 2022a), by directly incorporating embeddings from modality-specific encoders, we propose a paradigm shift for incorporating user history into LLMs. Instead of relying on computationally expensive textual representations, we treat user timelines as a distinct modality, directly embedding them for LLM contextualization.
This embedding-based approach offers several key advantages. First and most, it significantly reduces the context window length required for LLM processing, leading to substantial improvements in inference efficiency. Second, it allows LLMs to directly capture the latent patterns and temporal dynamics within user interactions, potentially fostering a deeper understanding of individual users and their evolving needs.
To realize this vision, we introduce User-LLM, a novel framework (1) that leverages user embeddings to directly contextualize LLMs. These embeddings, generated by a user encoder pretrained using self-supervised learning on diverse user interactions, encapsulate latent user behaviors, interests, and their evolution over time. We integrate these user embeddings with LLMs through cross-attention, enabling the models to dynamically adapt their responses based on the rich context of a user’s past actions and preferences.
By representing user timelines directly as embeddings, we drastically reduce the computational burden on the LLM during inference, leading to inference speedups of up to 78.1X compared to text-prompt-based methods. Our comprehensive experiments across three datasets further demonstrate User-LLM’s ability to outperform text-prompt-based contextualization on tasks requiring deep user understanding, particularly with long context inputs.
@@
  • We introduce User-LLM, a framework that leverages user embeddings to directly contextualize LLMs, enabling them to dynamically adapt to a user’s past actions and preferences. User-LLM supports various user encoder architectures and encoder-LLM co-training strategies.
  • We conduct extensive experiments on three public datasets, demonstrating that User-LLM significantly outperforms text-prompt-based contextualization methods, achieving up to 78.1X inference speedup and up to 16.33% improvement in performance on tasks requiring deep user understanding, particularly with long context inputs.
  • We provide a comprehensive analysis of the impact of different user embedding generation architectures, encoder-LLM co-training strategies, and different ways to integrate embedding with LLMs on LLM personalization, offering valuable insights for future research.
  • We delve into cross-attention mechanism for integrating user embeddings, exploring gated cross-attention to understand how user embeddings influence LLM behavior, and incorporate Perceiver layers for further efficiency gains.
@@ 2 Related Work
2.1 Multimodal LLM
Early work in multimodal LLMs primarily focuses on aligning images and text representations (e.g., CLIP (Radford et al. 2021), ImageBind (Girdhar et al. 2023)). Subsequently, fusion techniques leveraging cross-attention (e.g., Flamingo (Alayrac et al. 2022), Coca (Yu et al. 2022)) and soft-prompt (e.g., PaLI (Chen et al. 2022a), Palm-E (Driess et al. 2023)) emerged. Recent works, such as NextGPT (Wu et al. 2023b), OneLLM (Han et al. 2023), and Anymal (Moon et al. 2023), explored unified frameworks for diverse input modalities. Additionally, end-to-end training of multimodal models has gained attention, as seen in Gato (Reed et al. 2022) and Gemini (Gemini Team Google 2023). Inspired by these advances, User-LLM focuses on directly contextualizing LLM with embeddings for user understanding.
2.2 Language Model based Personalization
Recent research has extensively explored incorporating user interaction history as text prompts for LLM personalization and recommendation (Petrov and Macdonald 2023; Kang et al. 2023; Xu, Hua, and Zhang 2023; Liu et al. 2023a; Lyu et al. 2023; Ji et al. 2023; Li et al. 2023; Wu et al. 2023a). However, directly feeding the entire user history can be computationally expensive due to the resulting long context lengths. Our approach leverages user embeddings for efficient LLM contextualization, significantly reducing computational overhead. While a recent study (Doddapaneni et al. 2024) also explored LLM personalization with user embeddings, it focused on deriving embeddings from text-based user activity and integrating them via soft-prompt. In contrast, User-LLM utilizes efficient user activity tokens and sophisticated fusion techniques (e.g., cross-attention). We conduct a comprehensive evaluation across various datasets and tasks, demonstrating the effectiveness of User-LLM.
2.3 Long Context in LLMs
Long context is crucial for models to incorporate long-term user data. Existing approaches for LLMs include extending the context window (e.g., via positional encodings remapping (Chen et al. 2023a; Jin et al. 2024) or by exposing relevant additional contexts available to a subset of attention layers (Tworkowski et al. 2023)), distill information from the long context (e.g., by using a separate retrieval model (Xu et al. 2023), a memory bank (Wang et al. 2023), or training special tokens (Ge et al. 2023; Mu, Li, and Goodman 2023; Chevalier et al. 2023)), modified attention computation (Ding et al. 2023; Liu, Zaharia, and Abbeel 2023; Chen et al. 2023b), and linear attention frameworks (e.g., structured state space models (Gu et al. 2023; Gu and Dao 2023; Tay et al. 2022)). Our method tackles long context by using a single token to represent each user event, which is compatible with existing long context techniques.
@@ 2.4 User Modeling
Traditional approaches for user modeling and personalization have widely employed techniques like dual encoders, self-supervised learning (Yao et al. 2021; Xie et al. 2022; Chen et al. 2022b; Xia et al. 2023; Cai et al. 2023; Yang et al. 2023), and pretrained models (e.g., BERT (Devlin et al. 2018), Bert4Rec (Sun et al. 2019), U-Bert (Qiu et al. 2021))) to generate effective user representations (Covington, Adams, and Sargin 2016a, b; Sountsov and Sarawagi 2016; Volkovs, Yu, and Poutanen 2017; Chidambaram et al. 2018; Gillick, Presta, and Tomar 2018; Yanga et al. 2018; Ma et al. 2018; Yi et al. 2019; Gillick et al. 2019; Yang et al. 2020; Jiang et al. 2020; Ning et al. 2021). While these approaches provide valuable insights into user behavior, they primarily operate outside the realm of LLMs. In contrast, our work focuses specifically on enhancing LLM capabilities for personalization by directly contextualizing them with user representations. This allows us to explore the potential of LLMs as the foundation for powerful, adaptive AI agents. Therefore, direct comparisons with non-LLM based user modeling techniques are not within the scope of this work.
3 Methodology
This section introduces the User-LLM framework for contextualizing LLM with user embeddings (Fig.2).
3.1 Problem Statement
Given a set of users $\mathcal{U}$ and a set of items $\mathcal{V}$, each user $u\in\mathcal{U}$ has a timeline sequence $S_{u}=(v_{1}^{(u)},v_{2}^{(u)},...,v_{n_{u}}^{(u)})$ where each $v_{i}^{(u)}\in\mathcal{V}$ represents an item the user interacted with in chronological order. Each item $v$ is associated with a set of features $\mathcal{M}$. Each feature $m\in\mathcal{M}$ has a vocabulary $V_{m}$ that maps its values to integer IDs. For example, if items are movies, then the feature could be $\mathcal{M}=(name,\>genre,\>rating)$, where $V_{name}$, $V_{genre}$ map text values to IDs and $V_{rating}$ maps numerical ratings to IDs. Our goal is to develop a framework to contextualize a decode-only LLM with $S_{u}$, enabling it to generate personalized response $\mathit{A}_{u}$ for user $u$ given a query $\mathit{Q}$.
3.2 User-LLM
@@
As shown in Fig.2 (b), User-LLM contains an ID-based user encoder and a decoder-only LLM. The inference includes two-stages: (1) the user encoder generates user embeddings; (2) the embeddings are integrated into the LLM through cross-attention, providing the LLM additional context and guidance to generate personalized responses.
Figure 2: Overview of User-LLM. (a) Autoregressive Transformer encoder pretraining. (b)LLM contextualization with user embeddings. Features from a user’s interaction history are encoded by the Autoregressive User Encoder and then integrated into the language model via cross-attention.
User Embedding Generation
User-LLM leverages a Transformer-based encoder to generate user embeddings from ID-based interaction sequences.
Let us consider a scenario where each item has $M$ features ($|\mathcal{M}|=M$). Representing each item with feature IDs, an ID-based user timeline sequence contains $\mathit{L}$ items can be denoted as
$s_{u}=[(x_{1,m_{1}},...,x_{1,m_{M}}),...,(x_{L,m_{1}},x_{L,m_{M}})]$
where $x_{i,m_{j}}=V_{m_{j}}[m_{j}^{v_{i}}]$ denotes the ID of feature $m_{j}$ for item $v_{i}$.
Each element ID has its unique embedding representation. To obtain a unified representation for an item, we combine embeddings from different features into a single embedding. Let $E_{m_{j}}\in\mathbb{R}^{|V_{m_{j}}|\times d_{j}}$ be the embedding matrix associated with $V_{m_{j}}$, where $d_{j}$ is the embedding dimension for $m_{j}$. For item $v_{i}$, the fused embedding is $f_{i}=[E_{m_{1}}(x_{i,m_{1}});E_{m_{2}}(x_{i,m_{2}});...,E_{m_{M}}(x_{i,m_{M}})]$, where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation. To reduce the dimension of the fused embedding, we project $f_{i}$ to $f_{i}^{\prime}$ with a dimension of $d$.
Autoregressive User Encoder
As illustrated in Fig. 2 (a), an Autoregressive Transformer, denoted as $\mathcal{T}$, takes the fused embeddings $F^{\prime}=[f_{1}^{\prime},f_{2}^{\prime},...,f_{L}^{\prime}]$ into the Transformer decoder, and outputs embeddings $E_{s_{u}}=\mathcal{T}(F^{\prime})$, where $E_{s_{u}}\in\mathbb{R}^{L\times d}$. These embeddings are used for LLM contextualization in the next stage.
@@
LLM Contextualization with User Embedding
@@
User-LLM integrates user embeddings with LLMs using cross-attention (see Fig. 2 (b)). The output embeddings from the user encoder are down sampled by a projection layer, and then are cross-attended with intermediate text representations within the LLM, similar to how Flamingo (Alayrac et al. 2022) works.
Let $E_{su}$ denote the encoder output, and $CrossAttn(Q,K,V)$ is the cross-attention mechanism. Note that for cross-attention, $Q$ is LLM input, whereas $K$ and $V$ are compatible with the dimension of cross-input $E_{su}$. Let $O_{i}$ and $I_{i+1}$ denote the output of the $i$-th and the input of $(i+1)$-th self-attention layers, respectively. Our framework enables cross-attention as follows:
$O_{i}$ $=\text{LayerNorm}(O_{i})$ (1)
$O_{i}$ $=\text{CrossAttn}(O_{i},E_{s_{u}},E_{s_{u}})+O_{i}$ (2)
$I_{i+1}$ $=\text{FeedForward}(O_{i})$ (3)
In order to further investigate the cross-attention mechanisms between encoders and LLMs, we also implemented gated cross-attention in User-LLM by inserting gated cross-attention-only layers between LLM’s self-attention stack. Flamingo-style gating mechanism (Alayrac et al. 2022) was employed to control the contribution of cross-attention signal to LLM. That is, a scalar between 0 and 1 is introduced to equation (2):
$O_{i}$ $=\boldsymbol{\tanh(\alpha_{i})}\times\text{CrossAttn}(O_{i},E_{s_{u}},E_{s_{u}})+O_{i}$ (4)
where $\alpha_{i}$ is a trainable, zero-initialized scalar gating the $i$-th cross-attention layer. Using this approach, we enabled cross-attention between encoder output and transformer layers in the LLM with layer-specific cross-attention weights regulated by $\alpha$.
Note that our framework is adaptable to different integration mechanisms including prepending embeddings as soft-prompt (Lester, Al-Rfou, and Constant 2021).
3.3 Training Framework
To effectively extract information from user interaction data and teach LLM to understand the information encoded in the user embeddings, we propose a two-stage training framework: user-encoder pretraining and encoder-LLM finetuning.
User-encoder Pretraining
In the first stage, we pretrain the Autoregressive Transformer on user interaction sequence so that it learns how to extract information from the input data. As described in Section 3.2, the Autoregressive Transformer takes in a sequence of ID-based user interactions and outputs embeddings $E_{s_{u}}$. For pretraining, the decoder output is projected back to the original feature dimensions using feature-specific projection matrices $W_{m_{1}}\in\mathbb{R}^{d\times d_{1}},W_{m_{2}}\in\mathbb{R}^{d\times d_{2}},...,W_{m_{i}}\in\mathbb{R}^{d\times d_{i}}$: $h_{m_{j},i}=W_{m_{j}}e_{i}$, where $e_{i}\in E$ is the $i$-th embedding output from the Transformer.
Loss Calculation
The cross-entropy loss for each feature is calculated using the logits of the feature-specific projection:
$\mathcal{L}_{m_{j}}=-\frac{1}{L}\sum_{i=1}^{L}\log(\text{softmax}(h_{m_{j},i}))(f_{m_{j},i+1}^{\prime})$
where $f_{m_{j},i+1}^{\prime}$ is the projected fused embedding for feature $m_{j}$ at step $i+1$. The overall loss is the average of cross-entropy losses across features: $\mathcal{L}=\frac{1}{i}\sum_{j=1}^{i}\mathcal{L}_{m_{j}}$
Encoder-LLM Finetuning
In the second stage, we connected pretrained Autoregressive Transformer and a decode-only LLM with cross-attention and finetune the entire User-LLM model to align user embeddings and LLM’s text representations. The model input has two parts: an ID-based user interaction sequence as the input to the Autoregressive Transformer ($X_{\mathcal{T}}=s_{u}$) and a query as the input to the LLM ($X_{LLM}=\mathit{Q}$).